OpenStudy (anonymous):

how many grams of fluorine are contained in 5 molecules of BF3

OpenStudy (anonymous):

ok lets suppose that by fluorine you mean F^- then: FIRST METHOD: n=N/L n(BF3) = 5 / 6,022 * 10^23 mol-1 n(BF3) = 8,3 * 10^-24 mol to continue: n= m / M -> m= n*M n(F-) / n(BF3)= 3/1 = 3 n(F-) = 3 * n(BF3) n(F-) = 2,5 * 10^-23 mol and now: m(F-) = n(F-) * M(F-) m(F-) = 2,5*10^-23 mol * 67,8 gmol-1 m(F-) = 4,7 *10^-22 g SECOND METHOD: n=N/L n(BF3) = 5 / 6,022 * 10^23 mol-1 n(BF3) = 8,3 * 10^-24 mol m(F-) = n(BF3) * M(BF3) m(F-) = 4,7*10^-24 mol * [3*Ar(F) / M(BF3)] m(F-) = 4,7*10^-24 mol * [(3*19)/67,8] gmol-1 m(F-) = 4,7 * 10^-22 g here was used [3*Ar(F) / M(BF3)] which is sort of "equivalence unit" THIRD METHOD: task says that you have 5 molecules of BF3 and you need mass of F- so you just use common sense and multiply 5*3 = 15 ions of F- and now: n(F-) = N(F-) / L n(F-) = 15 / 6,022 * 10^23 mol-1 n(F-) = 2,5 * 10^-23 mol m(F-) = n(F-) * Ar(F-) m(F-) = 2,5 * 10^-23 mol * 19 gmol-1 m(F-) = 4,7 * 10^-22 g